OpenStudy (anonymous):

Factorize x^4-11x^2-1

jhonyy9 (jhonyy9):

sign x^2=y than you get y^2 -11y -1=0 y_1,_2=(11+/- sqrt(121+4))/2 =(11+/- sqrt125)/2 = (11+/- 5sqrt5)/2 x^2=(11+/- 5sqrt5)/2 x_1,_2=(11+/- 5sqrt5)/2 x_3,_4=-(11+/- 5sqrt5)/2